# MACE-OFF23: Transferable Machine Learning Force Fields for Organic Molecules

### v2

@@ III.6 Peptide Simulations
In the following section, we benchmark the performance of MACE-OFF23(S) on several well-studied large-scale biomolecular systems. We selected the small model since it is capable of accurately reproducing the properties of bulk water (Section III.4), while also having the computational speed required to simulate large systems over long timescales. This particular application area is of key interest, and state-of-the-art classical protein force fields have been carefully parameterized over many decades to reproduce key quantum mechanical and thermodynamic properties. However, this extensive parameterization process makes it difficult to extend classical force fields, and overall accuracy and transferability remain fundamentally limited by their functional form, which typically becomes apparent only when each generation of advancing hardware enables sufficiently long simulations to identify issues.
III.6.1 Ala3 Free Energy Surface
@@
We first studied the ability of MACE to reproduce the free energy surface of the Ala3 tripeptide. Using the MACE-OFF23(S) force field, we observed computational performance of 7.5 ns/day for the vacuum system, and 2.2 ns/day for the solvated system, using a 1 fs timestep. Although this remains significantly more expensive than the MM simulation (around 300 ns/day with a 2 fs timestep), it nevertheless enables sufficient sampling of the potential energy surface in a reasonable compute time.
@@
Figure 7 shows the free energy surface of Ala3 in water, modelled by AMBER14/TIP3P and MACE, respectively, sampled along the central $\phi$ and $\psi$ backbone angles. The AMBER14 simulations identify several local minima, corresponding to the anti-parallel $\beta$-sheet ($\phi<-120^{\circ},\psi>120^{\circ}$), a right-handed $\alpha$-helix ($\phi=-60^{\circ},\psi>-60^{\circ}$), the corresponding left-handed $\alpha$-helix ($\phi=60^{\circ},\psi>60^{\circ}$), and a polyproline II (PPII)-type structure ($\phi=-60^{\circ},\psi>120^{\circ}$). Compared to experimental fits of the NMR J-coupling constants [81], the AMBER14 simulation is generally in good agreement, with a slight overpopulation of both $\alpha$-helical structures and underpopulation of the PPII mesostate.
Figure 7: Free energy surfaces of Ala3 in explicit solvent. Free energies have been normalized to zero at the minimum to aid visual comparison.
[…]
III.6.2 Folding dynamics of Ala15
@@
Having confirmed that MACE is capable of constructing the free energy surface of a small peptide, we next investigated the folding of the longer helical peptide Ala15. Since this is a significantly larger system size than the dipeptide configurations that the model was trained on, this represents a nontrivial test of the extrapolation capability of the potential, including its ability to capture complex hydrogen bonding patterns required to stabilize the secondary structure. Such simulations are believed to be particularly difficult with purely local models such as MACE-OFF23.
As an initial test, the fully extended Ala15 structure was prepared and simulated in vacuo with MACE-OFF23(S). The top right panel of Figure 8 shows the assignment of the secondary structure during the simulation. The peptide rapidly folded into a stable helical structure, which was confirmed using hydrogen bonding counts.
We also measured the secondary structure of the peptide during trajectories collected at three higher temperatures. This analysis confirmed that the helical secondary structure is stable up to around 700 K, while at 800 K the hydrogen bond count significantly decreases, consistent with the experimental unfolding temperature of 725 K as shown in the two bottom panels of Figure 8.
Figure 8: Ala15 folding dynamics. Left: Hydrogen bond count using the DSSP algorithm [82] along the Ala15 trajectory at four different temperatures. Right: Secondary structure of Ala15 during the folding trajectory.
III.6.3 Protein simulation in explicit solvent
@@
Finally, we investigated the simulation of a fully solvated protein with MACE-OFF23. As a test case, we chose crambin, a 42 residue threonin storage protein containing four charged residues. This test goes well beyond what can typically be expected from a short-ranged force field trained on neutral species only.
@@
Using the MACE-OFF23(S) model, we first confirmed that the force field was indeed capable of simulating a large biomolecular system containing charged residues. In particular, it has recently been shown that the FENNIX potential fails to simulate a solvated 28 residue peptide due to unphysical proton transfers involving charged residues [37]. We confirmed that the protein backbone root-mean-square fluctuations relative to the minimized structure remained below 1 Å for the duration of a 125 ps simulation, no bond breaking occurred, and secondary structure motifs remained intact.
We then probed the ability of MACE-OFF23(S) to capture the key vibrational modes of the system. To this end, we prepared a simulation box containing crambin in the high hydration state, i.e. fully solvated in explicit water, for a total of approximately 18,000 atoms. The power spectrum was calculated from a 125 ps simulation recorded at 1 fs resolution. Experimentally, this system has been well studied, and key THz-region vibrational modes have been assigned [83]. This system has also been investigated with the AMBER classical force field and the GEMS machine learning potential [40, 84], where in the latter case a bespoke model was trained to simulate crambin by augmenting a training set of small molecules with large fragments of the solvated protein taken from classical MD simulation and evaluated using DFT.
Despite the lack of such system-specific “top-down” training data, Figure 9 shows that MACE-OFF23(S) is able to reproduce the key elements of the power spectrum, including the bending and stretching modes of water (1600 cm-1 and 3600 cm-1 respectively), further demonstrating the ability of MACE to accurately describe bulk water. In the low wavenumber region, the key internal motions of the protein and interactions with water are also captured. In particular Figure 9 shows a peak at 60 cm-1, which has been attributed to localized internal side chain fluctuations [83]. This peak is captured by all three force fields, with only small deviations in the peak position.
On the contrary, the prominent band located at around 220 cm-1 is clearly defined only in the GEMS and MACE simulations, whilst this appears as a shoulder in the AMBER simulation. This is thought to correspond to a stretching mode of the water molecules present in the solvation shell, and agrees with the experimental spectrum (inset) [83]. This is intriguing, as it suggests that both ML potentials, including the strictly local one presented here, are capable of reproducing complex interactions between the protein and structured solvation shells, whilst the AMBER classical force field does not predict a distinct peak here.
@@
The results of this sections III.5 and III.6 highlight the most significant difference between previous transferable organic ML potentials and the MACE-OFF23 models. These are, to our knowledge, the first ML potentials capable of reproducing experimental properties of condensed phase molecular systems, as well as large-scale, all-atom, explicit solvent biomolecular simulations without relying on augmented, system-specific training data.
@@
Figure 9: Protein in explicit solvent. Power spectrum of the high hydration state of crambin computed from 125 ps simulation. Inset shows correspondence of the MACE-OFF23(S) spectrum with the experimental spectrum for low wavenumber regions. GEMS and AMBER spectra are reproduced from Ref. [84].
@@ III.7 Computational performance
To evaluate the computational performance of MACE-OFF23, we perform NVT molecular dynamics for water using the LAMMPS software, with simulation sizes ranging from $10^{2}$ to $10^{5}$ atoms. The results are summarized in Figure 10; all simulations were performed with the MACE-OFF23(S) model and utilized NVIDIA A100 GPUs. The LAMMPS evaluator relies on LibTorch (the C++ interface for PyTorch) and the Kokkos performance portability libraries.
The black curve in Figure 10 shows the computational performance as a function of system size for single-GPU evaluation, where performance is measured in units of millions of timesteps per day (equivalent to ns/day if the timestep is 1 fs). The model achieved 6.5 ns/day peak performance for roughly 200 atoms and 2.5 ns/day for roughly 1000 atoms, which is competitive with other machine learning force fields such as ANI [27] or AIMNet [8].
The three colored curves demonstrate the performance when multiple GPUs are employed to access larger systems with domain decomposition. Each has a dashed part which connects the single-GPU result (on the black curve) to the two-GPU result; the notable decrease in performance is explained by a different treatment of “ghost” atoms (those atoms which lie outside a given subdomain, but sufficiently near the boundary to influence “local” atoms). Without domain decomposition (assuming periodic boundary conditions), each ghost atom is a translation of a local atom—or, put differently, the full structure comprises a periodic graph—allowing for efficient reuse of intermediate calculations and better performance. Once the full domain is split across GPUs, it becomes more challenging to reuse intermediate results efficiently and, at present, some redundant calculations are performed. (The isolated colored dots show the single-GPU performance without efficient handling of ghost atoms). Finally, the solid colored lines demonstrate perfect weak scaling, where each subdomain belongs to one GPU, and the simulation speed is now independent of the system size.
[…]
@@ IV Outlook
In this paper, we presented the series of MACE-OFF23 models, demonstrating the broad applicability, transferability and capability of purely short-range machine learning potentials for organic (bio)molecular simulations. We have shown that models, based on the MACE higher order equivariant message passing architecture, can improve on the pioneering ANI models [27] that served as the only widely applicable machine learning force fields for molecules for a number of years. We have demonstrated significant improvements both in terms of accuracy and extrapolation combined with high computational speed.
The lack of explicit long-range interactions limits the domain of applicability of the present model to neutral, non-radical, and non-reactive systems. This is something that the recently published AIMNet-2 model aims to address by extending the ANI models to include charged species and long-range interactions [8]. We are currently working on a next-generation MACE-OFF model that will similarly include an explicit description of charges, enabling the description of amino acids with different protonation states, charged nucleic acids, and counter-ions. This will pave the way towards obtaining an accurate quantum mechanical transferable machine learning force field for simulating the full range of biologically relevant systems.
@@ Acknowledgements
DPK acknowledges support from AstraZeneca and the EPSRC. DJC and JH acknowledge support from a UKRI Future Leaders Fellowship (grant MR/T019654/1). V.K. acknowledges support from the Ernest Oppenheimer Early Career Fellowship and the Sydney Harvey Junior Research Fellowship, Churchill College, University of Cambridge. W.C.W. acknowledges support from the EPSRC (Grant EP/V062654/1). JHM acknowledges support from an AstraZeneca Non-Clinical PhD studentship, and thanks Prof. Ola Engkvist, Dr Marco Klähn and Dr Graeme Robb from AstraZeneca for their helpful discussions and supervision. The authors also thank Christoph Schran for providing the carved water cluster geometries. We also thank Dr Peter Eastman and Dr Stephen Farr for their help implementing the MACE-OpenMM interface. We are grateful for computational support from the Swiss National Supercomputing Centre under project s1209, the UK national high-performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and the EPSRC grant ref EP/P022561/1, and the Cambridge Service for Data Driven Discovery (CSD3).
We used computational resources of the UK HPC service ARCHER2 via the UKCP consortium and funded by EPSRC grant EP/P022596/1. This work was also performed using resources provided by the Cambridge Service for Data Driven Discovery (CSD3). Access to CSD3 was also obtained through a University of Cambridge EPSRC Core Equipment Award (EP/X034712/1). The Rocket High Performance Computing service at Newcastle University was also used.
@@
DPK conceived the project, curated the training data, trained and evaluated the models, ran parts of the torsion drive tests, the molecular crystal tests, fitted the dipole and polarizability models, and wrote the initial draft. DPK and JHM implemented the OpenMM interface. JHM performed experiments on molecular liquids, peptides and proteins and wrote the initial draft. JTH implemented the MACE-torsion drive interfaces, and ran the DFT and force field torsion drives. VK ran the MD and PIMD simulations of water and molecular crystals and produced the IR and Raman spectra. IB and NJB implemented the accelerated custom kernels. WCW implemented the MACE-LAMMPS interface and benchmarked the model performance. IBM developed the intermolecular error analysis and evaluated the models. DJC and GC supervised the project. All authors discussed the results and edited the final manuscript.

### v3

@@ III.6 Biomolecular simulations
In the following section, we benchmark the performance of the MACE-OFF models on several well-studied large-scale biomolecular systems. This particular application area is of key interest, and state-of-the-art classical protein force fields have been carefully parameterized over many decades to reproduce key quantum mechanical and thermodynamic properties. However, this extensive parameterization process makes it difficult to extend classical force fields, and overall accuracy and transferability remain fundamentally limited by their functional form. This typically becomes apparent only when each generation of advancing hardware enables sufficiently long simulations to identify issues.
III.6.1 Extended cutoff MACE-OFF models
One key observation that emerged from validation of the MACE-OFF23(S) and MACE-OFF23(M) models was an overestimation of the density of liquid water by approximately 20% (Figure 9). Since there is extensive coverage of water clusters in the training set, we hypothesized that this resulted from a missing long-range contribution beyond the receptive field. To investigate this hypothesis, we fitted an additional model, MACE-OFF23(M)b, which has an increased 6 Å layerwise cutoff compared to 5 Å. The larger receptive field leads to a predicted water density within 2% of the experimental value at room temperature.
We also investigated the effect of including additional training data, made up of the Solvated Pubchem and Amino Acid Ligand Pairs datasets from SPICE v2 [63]. Given the direct applicability of these training data to biomolecular simulations, in the following section we test an additional model, MACE-OFF24(M), which includes the additional training data. Given the encouraging performance of MACE-OFF23(M)b, this model also employs the extended 6 Å cutoff.
Figure 9: Water density as a function of temperature for the extended cutoff MACE-OFF models
III.6.2 Ala3 Free Energy Surface
@@
We first investigated the ability of MACE-OFF24(M) to reproduce the free energy surface of the Ala3 tripeptide. We observed computational performance of $9.6\times 10^{6}$ steps/day for the vacuum system, and $2.2\times 10^{6}$ steps/day for the solvated system on an NVIDIA 80GB A100 GPU via the OpenMM interface. Although this remains significantly more expensive than the MM simulation (around $150\times 10^{6}$ steps/day), it nevertheless enables sufficient sampling of the free energy surface.
@@
Figure 10 shows the free energy surface of Ala3 in water, modelled by AMBER14/TIP3P and MACE-OFF24(M), respectively. Well-tempered metadynamics was used to enhance sampling along the central $\phi$ and $\psi$ backbone angles. The AMBER14 simulation identifies four local minima, corresponding to the anti-parallel $\beta$-sheet ($\phi<-120^{\circ},\psi>120^{\circ}$), a right-handed $\alpha$-helix ($\phi=-60^{\circ},\psi>-60^{\circ}$), the corresponding left-handed $\alpha$-helix ($\phi=60^{\circ},\psi>60^{\circ}$), and a polyproline II (PPII)-type structure ($\phi=-60^{\circ},\psi>120^{\circ}$). Comparing the dihedral angle distribution with NMR J-coupling constants, the AMBER14 simulation is generally in good agreement, with a slight overpopulation of both $\alpha$-helical structures and underpopulation of the PPII mesostate [85]. We therefore use this to enable visual comparison with a well-tuned classical forcefield.
The corresponding MACE simulations show a very similar free energy surface to AMBER14/TIP3P, with the same four local minima identified. Whilst the relative depths of the α-helical conformations are in agreement, AMBER predicts the free energy of the anti-parallel β-sheet to be 1 kcal/mol higher than the PPII minimum, compared to 0.1 kcal/mol for MACE.
We also compare the unbiased conformational ensemble predicted by MACE directly to experiment by calculating 3J coupling constants from the Ramachandran distribution. Computed values are shown in Table 3 in addition to literature values for AMBER14SB, ANI-2x and experiment [85, 86, 87]. MACE-OFF24(M) accurately predicts coupling constants, and is comparable with AMBER14SB. In agreement with our previous findings, both methods significantly outperform ANI-2x, highlighting the difficulties in computing accurate dynamics in the solution phase with MLPs.
3J(HN,Hα) 3J(HN,Cβ) 3J(HN,C)
MACE-OFF24(M) 5.70 1.64 1.37
ANI-2x 4.5 1.2 1.4
AMBER14SB 6.07 1.87 1.14
EXP 5.68 2.39 1.29
Table 3: Summary of 3J couplings computed from dihedral distributions. AMBER14SB and Experimental values taken from Ref. [85]. ANI-2x data taken from Ref. [86].
Figure 10: Free energy surfaces of Ala3 in explicit solvent. (left) Amber14SB/TIP3P and (right) MACE-OFF24(M).
III.6.3 Folding dynamics of Ala15
@@
Having confirmed that MACE-OFF24(M) is capable of recovering the free energy surface of a small peptide, we next investigated the folding of the longer helical peptide Ala15. Because this is a significantly larger system size than the dipeptide configurations on which the model was trained, this represents a nontrivial test of the extrapolation capability of the potential, including its ability to capture complex hydrogen bonding interactions required to stabilize the secondary structure. Such simulations are believed to be particularly difficult with purely local models such as those in the MACE-OFF family.
As an initial test, the fully extended Ala15 structure was simulated in vacuo. Figure 11 shows the assignment of the secondary structure during the simulation. The peptide folded within 200 ps, adopting initially a “wavy” intermediate structure, before first folding into a 310 helix. The secondary structure oscillates between the α-helix and the 310 helix for the remainder of the simulation, with the α-helix being the predominant motif. These results are in agreement with other works [9, 39], in which polyalanine peptides are observed to fold through a “wavy intermediate” and oscillate between 310 and α-helices. The oscillatory behaviour has also been observed experimentally [88].
To better understand the performance of MACE-OFF24(M), we compared the predictions to those of MACE-OFF23(M) and MACE-OFF23(M)b. Through ablation studies, we found that both the inclusion of additional training configurations and the increased radial cutoff were required to observe the oscillatory behavior, whilst all other models overstabilize the α-helix, and do not predict the 310 structure. We hypothesize that additional sampling of the nonbonded interactions present in the expanded training set, as well as a larger receptive field of 12 Å allows the model to learn more subtle hydrogen bonding interactions required to interconvert between the two helical motifs.
We also measured the secondary structure of the peptide as the temperature was increased. Starting from the folded structure at 300K, the temperature was increased in steps from 300K to 900K over the course of 3 ns (Figure 12). The folded structure remains stable up to around 500 K, after which point the helix unfolds and the structure is primarily a random coil. This is an underprediction compared to the experimental unfolding temperature of 725 K, and suggests that there are additional subtle stabilising interactions that are not fully captured by the model [89].
Figure 11: Ala15 folding dynamics. Secondary structure assignment during 500 ps trajectory generated with MACE-OFF24(M)FF24M. The structure proceeds via a “wavy” intermediate and oscillates between the α- and 310 helices. Secondary structure assignment was performed with the STRIDE algorithm [90].
Figure 12: Ala15 Temperature dependence. Secondary structure assignment with increasing temperature.
III.6.4 Protein simulation in explicit solvent
@@
Finally, we investigated the simulation of a fully solvated protein with MACE-OFF. As a test case, we chose Crambin, a 42 residue threonin storage protein that contains four charged residues. This test goes significantly beyond what can typically be expected from a local force field trained on neutral species only.
@@
Using the MACE-OFF24(M) model, we first confirmed that the force field was indeed capable of simulating a large biomolecular system containing charged residues. In particular, it has recently been shown that the FENNIX potential fails to simulate a solvated 28 residue peptide due to unphysical proton transfers involving charged residues [37]. We confirmed that the root mean square fluctuations of the protein backbone relative to the minimized structure remained below 1 Å for the duration of a 1 ns simulation, no bond breaking occurred and the secondary structure motifs remained intact (see SI Figure 17).
We then investigated the ability of MACE-OFF24(M) to capture the key vibrational modes of the system. To this end, we prepared a simulation box containing crambin in the high-hydration state, that is, fully solvated in explicit water, for a total system size of approximately 18,000 atoms. The power spectrum was calculated from the final 125 ps of a 1.6 ns simulation recorded at 2 fs resolution, to match the simulation length and sampling frequency employed in Ref [41]. Experimentally, this system has been well studied, and key THz region vibrational modes have been assigned [91]. This system has also been investigated with the AMBER classical force field and several machine learned potentials, including GEMS and, more recently, SO3LR [40, 39, 9]. All three MLPs produce spectra in qualitative agreement with each other. The water peaks at 1640 cm-1 and 3200-3600 cm-1 are reproduced reasonably accurately, especially considering that no nuclear quantum effects are taken into account (Figure 13). Furthermore, in the terahertz region, two peaks are identified by both GEMS, SO3LR and MACE-OFF24(M). The most obvious of these occurs at 60 cm-1 and corresponds to localised internal side chain fluctuations, while a peak at 220 cm-1 is due to specific configurations of water at the protein surface.
We note that, whilst a 125 ps trajectory is sufficient to qualitatively identify the spectral features, a much longer simulation of at least several nanoseconds would be required to fully converge the higher frequency THz region of the spectrum and enable full quantitative comparison with experiment, as demonstrated for the GEMS model in Ref. [39]. We note that the power spectrum computed on the final 1.3 ns simulation (discarding the initial 300 ps to equilibration) does not yield a significantly different spectrum to that of the 125 ps simulation used in Figure 13, indicating that significantly more sampling would be required to resolve the THz lineshapes sufficiently to enable experimental comparison.
It is also noteworthy that whilst the three MLPs compared in this section give a qualitatively similar result, their modelling choices are significantly different. Most notably, GEMS and SO3LR explicitly include long-range interactions, whilst MACE is local. Additionally, of these models, both MACE and SO3LR are universal potentials, whilst GEMS is fitted in a system-specific fashion. We mention that, in addition to the GEMS model, the authors benchmarked a variant which was trained only on small fragments, denoted GEMS*, instead of the large system-specific configurations used to fit GEMS [39]. The power spectrum produced by this model is also in qualitative agreement with the other ML potentials, however, the RMSD is reduced, indicating the protein structure is confined to small fluctuations around local minima. This model also fails to reproduce experimental observables of other biomolecular systems, for example Ala15 is not predicted to fold under the GEMS* potential, suggesting the top-down fragments are required to accurately reproduce nonbonded interactions within the GEMS family of models.
@@
The results of sections III.5 and III.6 highlight the most significant differences between previous transferable organic ML potentials and the MACE-OFF models. These are, to our knowledge, the first local ML potentials capable of reproducing experimental properties of condensed phase molecular systems, as well as large-scale, all-atom, explicit solvent biomolecular simulations without relying on augmented, system-specific training data.
@@
Figure 13: Protein in explicit solvent. Power spectrum of the high hydration state of crambin computed from the final 125 ps of a 1.6 ns simulation with MACE-OFF24(M). GEMS, AMBER and SO3LR spectra are reproduced from Ref. [41].
@@ III.7 Computational performance
To evaluate the computational performance of the MACE-OFF models, we performed NVT molecular dynamics for water using both OpenMM and LAMMPS, with simulation boxes ranging from $10^{2}$ to $10^{5}$ atoms. The OpenMM results shown in Figure 14 were collected with a single NVIDIA A100 80GB GPU. The LAMMPS results shown in Figures 15-16 highlight both CPU and GPU evaluation.
In addition to the native PyTorch implementations of the MACE-OFF23 models, we also report results for optimized implementations leveraging custom C++, CUDA, and/or Kokkos code. One optimized variant, MACE-OFF23(S)-OPT, incorporates custom CUDA kernels which are packaged as a standalone, inference-only library, cuda_mace [92]. Another pure-C++ implementation, as well as a Kokkos implementation, will be made available as part of the symmetrix library [93].
In these optimized variants, the learnable radial function $R^{(t)}_{k\eta_{1}l_{1}l_{2}l_{3}}(r_{ij})$ is evaluated with cubic splines, which efficiently map $r_{ij}$ to each of the outputs of the radial MLP. Furthermore, Eqs. II.1 and 3 are implemented in unified kernels that simultaneously compute the one-particle basis and perform the summation over neighbors to generate the invariant atomic basis $A_{i}$. The computation of spherical harmonics, as detailed in equation II.1 is accelerated using the sphericart [94] library, which offers a highly optimized implementation for real spherical harmonics. All linear and equivariant linear operations have been substituted with custom kernels, and the cuda_mace implementation further leverages tensor core acceleration while incorporating error-correction techniques to ensure numerical accuracy. Additionally, the computation of the product basis $\mathbf{B}_{i}$ in equation 4 is optimized using the sparsity of the generalized Clebsch-Gordon coefficients $\mathcal{C}^{LM}_{\eta_{\nu}\bm{l}\bm{m}}$, significantly accelerating the product calculations over both the body order $v$ and the coupling terms.
Figure 14: Performance in OpenMM. Performance of MACE-OFF23 in OpenMM for liquid water in the NVT ensemble at 1 g/cm3 at 300K with a 1 fs timestep on a single NVIDIA 80GB A100 GPU.
Figure 15: Throughput in LAMMPS. Performance of molecular dynamics in LAMMPS for water in the NVT ensemble at 1 g/cm3 and 300 K with a 1 fs timestep. All simulations used variants of the MACE-OFF23(S) model.
The left panel highlights CPU performance achievable on the UK-based Archer2 CPU supercomputer, using the new pure-C++ MACE implementation. The solid black curve demonstrates the throughput of a single 128-core node as a function of the number of atoms. The two dotted curves demonstrate weak scaling with 432 atoms/node or 3600 atoms/node. All calculations used double precision floating point numbers.
The right panel demonstrates the performance of the GPU evaluators. The three solid curves provide single-GPU results (determined with NVIDIA A100 80GB GPUs) for the native Torch-based MACE-OFF23(S), a custom Kokkos-based MACE implementation, and the cuda_mace implementation marked as OPT*. The asterisk indicates that these latter simulations used single precision floating point numbers, with error correction for the most numerically sensitive terms, in contrast to the remaining calculations which were performed with double precision. Finally, the two dashed curves demonstrate the performance when multiple GPUs are employed, using the pure-Kokkos implementation, to access larger systems with domain decomposition. More work is needed to improve this weak scaling.
Figure 16: Evaluation speed in LAMMPS. Data from Figure 15, replotted to show the average duration of each timestep.
@@ IV Outlook
In this paper, we presented the series of MACE-OFF force field and dipole models, demonstrating the broad applicability, transferability and capability of purely short-range machine learning potentials for organic (bio)molecular simulations. We have shown that the models, based on the MACE higher order equivariant message passing architecture, can improve on the pioneering ANI models [27] that served as the only widely applicable machine learning force fields for molecules for a number of years. While a full ablation study between different ML models is not possible due to a wide range of design choices made in their construction, the MACE-OFF model architecture, training data and loss function all contribute to significant improvements in both accuracy and extrapolation, combined with high computational speed.
The lack of explicit long-range interactions limits the domain of applicability of the present model to neutral, non-radical and non-reactive systems. This is something that the recently published AIMNet2 model aims to address by extending the ANI models to include charged species and long-range interactions [8]. We are currently working on a next-generation MACE-OFF model that will similarly include an explicit description of charges, enabling the description of amino acids with different protonation states, charged nucleic acids, and counter-ions. This will pave the way towards obtaining an accurate quantum mechanical transferable machine learning force field for simulating the full range of biologically relevant systems.
@@ Acknowledgements
DPK acknowledges support from AstraZeneca and the EPSRC. DJC and JH acknowledge support from a UKRI Future Leaders Fellowship (grant MR/T019654/1). V.K. acknowledges support from the Ernest Oppenheimer Early Career Fellowship, the Sydney Harvey Junior Research Fellowship, Churchill College, University of Cambridge, and UCL’s startup funds. W.C.W. acknowledges support from the EPSRC (Grant EP/V062654/1), as well as an engagement from the Virtual Institute for Scientific Software at Georgia Tech, funded by Schmidt Sciences. In particular, Ketan Bhardwaj and Dave Brownwell of the Georgia Tech Scientific Software Engineering Center contributed to the Kokkos-based MACE evaluator. JHM acknowledges support from an AstraZeneca Non-Clinical PhD studentship, and thanks Prof. Ola Engkvist, Dr Marco Klähn and Dr Graeme Robb from AstraZeneca for their helpful discussions and supervision. The authors also thank Christoph Schran for providing the carved water cluster geometries. We also thank Dr Peter Eastman and Dr Stephen Farr for their help implementing the MACE-OpenMM interface.
We are grateful for computational support from the Swiss National Supercomputing Centre under project s1209, UCL Myriad High Performance Computing Facility, the Rocket High Performance Computing service at Newcastle University, the UK national high-performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and EPSRC grant EP/P022561/1, the Cambridge Service for Data Driven Discovery (CSD3), obtained through a University of Cambridge EPSRC Core Equipment Award (EP/X034712/1), and Ångström AI.
@@
DPK conceived the project, curated the training data, trained and evaluated the models, ran parts of the torsion drive tests, the molecular crystal tests, fitted the dipole and polarisability models for water and paracetamol bulk systems, and wrote the initial draft. JHM and DPK implemented the OpenMM interface. JHM trained and evaluated the models, curated the datasets, performed experiments on molecular liquids, peptides and proteins, benchmarked the model performance in OpenMM and wrote the initial draft. JTH implemented the MACE-torsion drive interfaces, and ran the DFT and force field torsion drives. VK fitted the MACE-OFF23-$\mu$, performed the MD and PIMD simulations of water and molecular crystals and produced their IR and Raman spectra. YP performed the error analysis for the MACE-OFF23-$\mu$ model and performed MD and PIMD simulations and calculated the IR spectrum of paracetamol molecule. IB and NJB implemented the accelerated custom kernels, and the cuda_mace library is maintained by NJB. WCW implemented the MACE-LAMMPS interface, wrote the pure C++ and Kokkos implementations, and benchmarked the model performance in LAMMPS. IBM developed the intermolecular error analysis and evaluated the models. DJC and GC supervised the project. All authors discussed the results and edited the final manuscript.
Data Availability
The data used to train the models are publicly available at: https://doi.org/10.17863/CAM.107498. The torsion drive dataset is also available at: https://zenodo.org/records/11385284. The MACE-OFF series of models are available at: https://github.com/ACEsuit/mace-off/tree/main and the MACE-OFF23-$\mu$ model is available at: https://github.com/venkatkapil24/MACE-OFF-mu-models